# Entanglement Entropy at Large-N

### v1

@@ Abstract
I show that at early times the evaporation process for a stack of ND5-branes at high energy is suppressed in the large-N limit. At much later times, the new saddles in the gravitational action are no longer suppressed at large-N, and evaporation proceeds as usual.
@@ 1 Motivations
The theoretical framework is defined on the world-volume of NS5-branes in the limit of vanishing string coupling, $\alpha^{\prime}\to 0\,.$ It was shown in [1] that LST is holographically dual to string theory on CGHS black hole. We shall rely heavily on this fact.
@@
The aim of this work is to clarify whether this pattern is due to the setup temperature or to some deeper cause. To do this, I look at the entanglement entropy in the spirit of [5]. Our results are unambiguous: the large-N limit suppresses semi-classical configurations, and as a consequence the system does not interact with the environment. Corrections, i.e. non-trivial saddles of the Euclidean path integral [6, 7], are not suppressed in this limit and are fully responsible for the unitary evaporation of the black hole.
For those familiar with QCD, this effect has a close parallel: the the decay $\rho\to\pi\pi$ at large-Nc [8].
@@ 3.1 Entanglement entropy for a single interval
Figure 1: Conformal diagram either for (2.2) or (2.7). Red curves depict the singularity, dashed lines the horizon and full black lines the asymptotic. We depict the configuration containing an island $\mathrm{I}$. The aim is to compute the entanglement entropy of the union $\mathrm{I}\cup\chi^{+}$.
@@ 3.2 Entaglement entropy for two-disjoint intervals
These islands can be located either inside or outside the horizon depending on the specific gravitational model. This is equivalent to look for the entanglement entropy of two disjoint intervals inside $\mathbb{R}^{2}$ [23] between the points $(y^{+}_{1},y^{-}_{1})$ and $(y^{+}_{2},y^{-}_{2})$ in vacuum state. The contribution of the matter can be factorised as followsº
@@ 3.2.1 $\mathcal{O}(\mathrm{c_{\mathrm{LST}}}(N))\lesssim\mathcal{O}(t)$
which is an extreme for the values

### v2

@@ Abstract
I show that at early times the evaporation process for a stack of NS5-branes at high energy is suppressed in the large-N limit. At much later times, the new saddles in the gravitational action are no longer suppressed at large-N, and evaporation proceeds as usual. This effect is due to the non-invariance of the metric under coordinate parameterisation. This fact introduces a parametric dependence in the thermodynamic quantities, which can lead to quantities that are suppressed in a particular corner of the parameter space.
@@ 1 Motivations
The theoretical framework is defined on the world-volume of NS5-branes in the limit of vanishing string coupling, $g_{\alpha}\to 0\,.$ It was shown in [1] that LST is holographically dual to string theory on CGHS black hole. We shall rely heavily on this fact.
@@
The aim of this work is to clarify whether this pattern is due to the setup temperature or to some deeper cause. To do this, I look at the entanglement entropy in the spirit of [5]. Our results are unambiguous: the large-N limit suppresses semi-classical configurations, and as a consequence the system does not interact with the environment. Corrections, i.e. non-trivial saddles of the Euclidean path integral [6, 7], are not suppressed in this limit and are fully responsible for the unitary evaporation of the black hole. In the spirit of [8] this new configurations will be complex solutions of the gravitational equations and correspond to saddles relevant to the unitary Page curve.
For those familiar with QCD, this effect has a close parallel in the decay $\rho\to\pi\pi$ at large-Nc. Inserting the hadronic spectrum of large–Nc QCD as a proxy for the real hadronic spectrum provides rather good approximation. In such a model the spectrum of the theory in the large–Nc limit consists of an infinite number of narrow stable meson states that cannot decay because of the zero width [9].
@@ 2. NS5
In Fig. (1) we represent the conformal diagram, $\{U,V\}$ coordinates (2.11), for (2.9).
Figure 1: Conformal diagram for either (2.2) or (2.7). Each point in the diagram represents a two-dimensional Euclidean space, in our case extending in the directions $\{x_{1},u\}$ of (2.9). Although the conformal factors (2.15, 2.16) are different, their causal structures are the same. Red curves show the singularity at $r\to 0$. Dashed lines represent the horizon, $r\to r_{0}$. Finally, solid black lines represent the asymptotic. One salient point is that the $r\to 0$ surfaces bend the conformal diagram, as does the AdS Schwarzschild black hole [21]. The region $\chi$, with states identified with the Haking radiation, is divided into two parts, $\chi^{-}$ and $\chi^{+}$. The boundary surfaces of $\chi^{\pm}$ are $y_{2}^{\pm}$ respectively. We show the configuration containing an island, $\mathrm{I}$, that extends outside the horizon. Its boundaries are located at $y_{1}^{\pm}$. The goal is to compute the entanglement entropy of the union $\mathrm{I}\cup\chi^{+}$.
@@ 3.2 Entaglement entropy for two-disjoint intervals
These islands can be located either inside or outside the horizon depending on the specific gravitational model. This is equivalent to look for the entanglement entropy of two disjoint intervals inside $\mathbb{R}^{2}$ [25] between the points $(y^{+}_{1},y^{-}_{1})$ and $(y^{+}_{2},y^{-}_{2})$ in vacuum state. The contribution of the matter can be factorised as follows
@@ 3.2.1 $\mathcal{O}(\mathrm{c_{\mathrm{LST}}}(N))\lesssim\mathcal{O}(t)$
which has an extreme for the values
@@ 5 Some Remarks
6 Some generalizations
It is tempting to ascribe the behaviour of (3.26) to the existence of a maximum temperature, but as we shall see, this behaviour seems to be inherent to the large-N limit and is most likely common to a larger family of gravitational backgrounds, namely those that are not scale invariant as described in section 4. Below we show one of such background.
The metric (2.7) is the ultraviolet completion of a large family group of regular non- abelian monopole solutions in ${\cal N}=4$ gauged supergravity, interpreted as 5-branes wrapped on a shrinking $S^{2}$ [29]. In the following we shall deal with a thermal deformation of one of such metrics dual to ${\cal N}=1$ SQCD with a superpotential coupled to adjoint matter [30].
The metric field in Einstein frameIn accordance with [26] we find that the value for the entanglement entropy is independent of the frame we have used. is given by
$ds^{2}=e^{\phi_{0}/2}r$$\left\{-K(r)dx_{1}^{2}+\sum_{j=2}^{4}dx_{j}^{2}+N\alpha^{\prime}\left(\frac{4}{r^{2}K(r)}dr^{2}+\frac{1}{\xi}d\Omega_{2}^{2}+\frac{1}{4-\xi}d\tilde{\Omega}_{2}^{2}\right)\right.$ (6.43)
$\left.+\frac{N\alpha^{\prime}}{4}\left(d\psi+\cos\theta d\varphi+\cos\tilde{\theta}d\tilde{\varphi}\right)^{2}\right\}\,,\quad K(r)=1-\left(\frac{r_{0}}{r}\right)^{4}\,.$
In addition we have a dilaton field which is linear, $\phi=\phi_{0}+r\,,$  and a RR 3-form field.
For this model the expression for (2.12) is almost identical to that obtained in the LST model (2.15)
$F_{\mathrm{}}(y)=\frac{b}{2}\log(y^{4}-1)\,,\quad c=\frac{1}{b}\,,\quad\Omega^{2}(y)=b^{2}\frac{r_{0}^{2}}{y^{3}}\,.$ (6.44)
After a bit of algebra one recover the suppressed behaviour at large-N, (3.26)
$S(\chi)=S^{b}_{\rm matter}$ $\approx\frac{c}{3}\frac{1}{\sqrt{N}}\frac{t_{y}}{l_{s}}\,,$ (6.45)
while, as expected, the corrections do not follow this pattern, as in (3.37).
So the main message to take away is that some gravitational models may not be invariant under coordinate reparametrisation. This fact introduces a parametric dependence in the thermodynamic quantities, which can lead to quantities that are suppressed in a particular corner of the parameter space.